# Information Discovery in e-Commerce

### v1

@@ 3.1.1 Basic concepts
An interface refers to an interactive component of a webpage or an application [hearst-2009-search]. Referring to all interactive components on e-commerce portals (e.g., search bars, navigation panels, lists of recommended items, item titles, and user reviews), e-commerce interfaces are playing invaluable role for the e-commerce user experience. Therefore, e-commerce interfaces dramatically impact the performance of an e-commerce platform. Depending on the nature of the stakeholders involved, most e-commerce sites can be divided into four types of business: business-to-business, business-to-consumer, consumer-to-consumer, and consumer-to-business [nemat2011taking]:
@@
B2C indicates scenarios where businesses directly sell items to consumers. Most online shopping platforms, such as Amazon, Booking.com, and JD.com belong to this type of business. Fig. 3.1(2) shows an item page from Amazon as an example of the interface used by B2C businesses.
@@ 3.1.2.1 Search results in e-commerce
E-commerce search engines provide category options with the search bar. During the early development of e-commerce search, interfaces of different types have been considered, e.g., devoted type, divided type, co-existing type, and multi-page type [lu2006clustering]. But with the development of e-commerce search, these types of interfaces have been blended by e-commerce platforms. Currently, a typical e-commerce search system includes three main components: query processing, candidate retrieval and ranking [zhang2020towards]. In query processing, the search engine rewrites a query from the user into a term-based representation that can be processed by downstream components. In the candidate retrieval stage, the system uses the inverted index to retrieve candidate products to match queries. Finally, the ranking component orders the retrieved candidates based on factors such as relevance, and predicted conversion ratio. We will discuss research about principles and strategies of all these three components in Chapter 5 with more details.
@@ 3.1.2.2 Recommendation results in e-commerce platforms
For many e-commerce platforms, recommendation is becoming the most important service to help users find their needed items. E.g., recommendations have been reported to contribute to the majority of both revenue and traffic in Taobao [wang2018billion], where one billion users can be connected to two billion items. To this end, the homepage on the mobile Taobao app is generated based on consumers’ past behavior via recommendation algorithms, as illustrated in Fig. 3.3.
@@
Different types of recommendation results may be shown at different stages of a customer’s. Examples include “substitutes” (see Fig. 3.4(a) and “complementary items” before and after the user adds a product to their cart (see Fig. 3.4(b) and 3.4(c), respectively). Once the consumer clicks a recommended product, the system will automatically jump to the product detail page, which includes product titles, product descriptions, categories, ratings, and reviews. We will discuss more details about strategies and technologies of e-commerce recommendation in Chapter 6.
@@ 3.1.2.3 Product titles in e-commerce platforms
Item title compression, also called short title extraction [gong2019automatic], is meant to extract sufficient words from lengthy and verbose titles to produce a succinct new title to improve the user experience on mobile devices [wang2018multi]. Inspired by neural extractive document summarization methods [ren2017leveraging], item title compression methods apply neural networks to weight the importance of each word in the item title. gong2019automatic proposed a feature-enriched neural extractive model to extract short titles. Specifically, the authors apply a recurrent neural network as a sequential classifier with three types of features: content, attention, and semantics respectively. By using user search logs as external knowledge, wang2018multi proposed a multi-task learning approach for improving item title compression. The proposed method is composed of two seq2seq components which share an identical encoder. The authors combine these two components with an overall pointer neural networks [vinyals2015pointer] to automatically select the most informative words from the given item title.
Pointer neural networks easily omit key information. To tackle this problem, sun2018multi proposed a multi-source pointer network model, named the multi-source pointer network (MS-Pointer), by considering two extra constraints: (i) irrelevant information reduction; and (ii) the key information retainment. Fig. 3.6 provides an overview of MS-Pointer, with two encoders. In MS-Pointer, in addition to the encoder for the source title, the authors add another knowledge encoder that uses an LSTM to embed the brand name and the commodity name. As shown in Fig. 3.6, MS-Pointer combines the original title “Nintendo switch console…” and background knowledge “brand name: Nintendo”, and then it generates the short title about the item “Nintendo switch”. Recently, DBLP:conf/emnlp/FetahuCRM23 proposed a instruction fine-tuning strategy to summarize product titles according to various criteria such as number of words in a summary or inclusion of specific phrases.
@@
The task of title generation has been proposed to extend the task of title compression into a text generation problem. Unlike title compression, which only extracts words from item titles, the task of title generation is to generate a short item title so as to address the problem of inaccurate item titles in e-commerce [zhang2019multi]. To generate a succinct and accurate short title from a long source title, zhang2019multi proposed a multi-modal generative adversarial network, named MM-GAN, which addresses the title generation task as a reinforcement learning problem. MM-GAN is composed of two main components, a title generator and a discriminator (Fig. 3.7). Given the source title and its corresponding tags or features, the generator applies an LSTM-based network to generate a short item title. The discriminator, i.e., a binary classifier, distinguishes whether the generated short titles are human-generated or machine-generated. Thus, an adversarial learning procedure is constructed, in which the quality of the short title depends on its ability to fool the discriminator into believing it is a human-generated one, and the output of the discriminator is a reward for the generator to improve the generation performance. Recently, scene marketing has become a new marketing mode for product promotion where scene scenarios are created to demonstrate product functions [zhao2020data]. To help the e-commerce system find scene topics, lin2022automatic proposed a topic generation method to generate scene-based titles in e-commerce.
@@ 3.1.2.4 Product descriptions in e-commerce platforms
As shown in Fig. 3.8, many e-commerce platforms provide a short description for each item so as to showcase the features of the item. As an important factor in content marketing, the item description is key for increasing consumer engagement. During the early years of e-commerce, item descriptions were usually written or edited by human copywriters. However, the availability of an increasing number of items in e-commerce makes this manual process too costly. Moreover, with the development of virtual assistants in e-commerce, such as Alexa and Tmall Genie, there is a growing demand for automatically generating a short description given item attributes. To address this demand, the task of item description generation has been proposed. Item description generation needs to generate an item’s description from a series of complicated attributes. wang2017statistical proposed a statistical framework to weight the relative importance of the attributes of an item and to maintain accuracy at the same time. In Fig. 3.9 we specify the framework of the proposed item description model. By combining sentence-level templates extracted from the input data with knowledge from a pre-trained dataset, the authors generate and rank candidate item descriptions through an online document planning stage.
@@
Unlike early studies that focused on generating item descriptions purely from the item’s attributes, zhang2019automatic generated a pattern-controlled item description from multiple features, e.g., titles and item categories. Based on the copy mechanism [gu2016incorporating], the authors propose a pattern-controlled pointer-generator network (PGPCN) to generate the description. In PGPCN, a transformer is applied in the encoder component, whereas the decoder is used to control the pattern (e.g., the category, the length, and the style of the description) of the item.
It is important that the descriptions generated for item description are grounded in facts. To generate a fact-based description, chanstick2019 proposed an encoder-decoder framework, called the fidelity-oriented product description generator (FPDG), by searching key information from keyword labels. The authors established semantic connections between item keywords and the generated product description. As shown in Fig. 3.10, FPDG has two main components: (i) a keyword encoder that stores the word and its entity label in the token memory and self-attention modules, and (ii) an entity-based generator that generates an item description based on the memory and self-attention modules.
@@
Personal interest is neglected by all of the above approaches that generate descriptions given attributes or keywords. To address this shortcoming, chen2019towards proposed a knowledge-based personalized item description generation strategy. The authors extended the encoder-decoder framework [sutskever2014sequence] to a sequence modeling formulation using a self-attention mechanism. A large variety of item attributes, including the target user’s personalized preference features, are combined in an attribute fusion component through multi-layer attention mechanisms; retrieved external knowledge is incorporated in a knowledge incorporation component. In Fig. 3.11, we provide an example of the knowledge-based personalized item description generation.
@@ 3.1.2.5 Question answering in e-commerce
Question-answering (QA) systems are designed to provide direct and concise answers to user queries by understanding the intent behind a question and retrieving or generating the most relevant information. QA systems aim to deliver specific answers from structured or unstructured data sources, such as web documents or knowledge bases. QA systems are essential across a variety of domains, including web search, customer support, where users seek quick, accurate, and contextually relevant information [radev2002evaluating, tapeh2008knowledge, yin2015neural]. To increase the number of sales, most e-commerce portals provide a QA service to facilitate the customers’ shopping procedure by answering their questions about products [gao2019product, feng2021multi]. Currently, on many e-commerce sites, a user can ask a question about a product, and the QA system allows some users (e.g., customers who bought this product) to provide answers [gao2019product]. In Fig. 3.12, we show examples of question-answering services at Amazon and JD.com. More detailed discussions of e-commerce QA is provided in Section 7.1.
@@ 3.1.2.6 User reviews in e-commerce
The sentiment classification task is to label a given text with a specific opinion label. It has received lots of attention during the past two decades [pang2002thumbs, go2009twitter, pan2010cross, kamal2012mining, tang2014learning, pontiki2015semeval, tsytsarau2016managing, sun2017review]. Work on sentiment classification in an e-commerce context has attempted to capture a user’s opinion about a specific item from reviews on an e-commerce platform [sun2017review, tang2015learning, xia2015dual, tripathy2016classification, li2018effect]. Traditional approaches to sentiment classification focus on the classification problem given textual attributes of an item [pang2002thumbs], while largely ignoring the relation between users and the item. To address this problem, tang2015learning proposed a neural network, the user-product neural network (UPNN), to incorporate user and item information into a document-level sentiment classification procedure. In particular, the authors jointly embedded the user preference information, i.e., ratings and reviews, and item attributes. Then, a convolutional neural network is used to predict the sentiment label of the target review.
Sentiment classification using the relation between users and products or services faces two important challenges: (i) the sparseness of user-item interactions, and (ii) the information in user embedding methods. chen2016neural proposed a fine-grained hierarchical neural network model to incorporate global user and item information into sentiment classification. Unlike many sentiment classifiers which use convolutional neural networks, the authors applied a hierarchical LSTM [gers1999learning] to jointly generate sentence-level representations and document-level representations. Then, user-item interaction information is applied as attention over various regions of a document to enhance the sentiment classification. wu2018improving distinguished between different roles of words and sentences in user reviews: to describe the user’s preferences or to describe an item’s characteristics. To distinguish between these roles, the authors proposed an attention-based hierarchical neural network model to embed user and item information to generate two text representations with user attention or item attention, respectively. Hao2021wwwsentiment leveraged fine-grained latent opinion knowledge into the sentiment classification process by using a variational reasoning method.
@@
Given the fact that an item can be commented on by hundreds of thousands of consumers, the quality of reviews in e-commerce varies considerably and not all reviews are helpful. To gain insights from helpful reviews, the task of review helpfulness prediction has attracted attention from both academia and industry [mcauley2016addressing]. Early studies on review helpfulness prediction employed feature-aware methods, where multiple types of features, such as structural features [kim2006automatically, susan2010makes, xiong2011automatically], emotional features [martin2014prediction], semantic features  [yang2015semantic], argument features [liu2017using], and lexical features [xiong2014empirical], were successfully applied.
Motivated by the progress of deep neural networks, fan2018multi introduced a multi-task neural learning (MTNL) architecture for identifying helpful reviews. chen2018cross proposed a CNN-based neural network with multi-granularity (i.e., character-level, word-level, and topic-level) features for helpfulness prediction. fan2019product proposed an end-to-end deep neural architecture to capture the intrinsic relationship between the meta-data of an item and its numerous comments that could be beneficial to discover the helpful reviews. Multi-modal data has become increasingly popular in online reviews. To analyze multi-modal reviews, liumulti2021 proposed a multi-modal review helpfulness prediction task that is aimed at exploring multi-modal clues for review helpfulness prediction. The authors proposed an item-review coherent reasoning module to capture the intra- and inter-modal coherence between the target item and the review. han2022sancl proposed a selective attention approach, including probe mask generation and mask-based attention computation, for the multi-modal review helpfulness prediction problem. To mine the mutual information of cross-modal relations in the input, nguyen2022adaptive proposed an adaptive cross-modal contrastive learning mechanism, with a multi-modal interaction module to correlate modalities’ features.
@@
Feature-aware methods are inspired by previous document summarization methods: yang2010feature proposed a feature-based item review summarization method to satisfy the detailed information needs of customers. To jointly summarize reviews and predict ratings in a mobile environment, liu2012movie proposed a latent semantic indexing based approach to extract features and attributes from user reviews and ratings. For new items without reviews, a probabilistic retrieval method is proposed to extract relevant opinion features from other items to describe the item information [park2015retrieval]. Another important part of review summarization concerns aspect extraction from user reviews [chen2014aspect, angelidis2018summarizing, bravzinskas2020few], where target entities and aspects need to be extracted from opinionated text. chen2014aspect extracted prior knowledge automatically from user reviews and propose a fault-tolerant model to extract aspects guided by the knowledge. Category hierarchy information is combined with a topic model to improve the performance of aspect extraction [yang2017aspect]. By jointly considering fine-grained aspect-topic-sentiment connections, tan2017sentence propose a generative topic aspect sentiment model.
With the development of deep learning, item review summarization has been tackled from a range of perspectives [ly2011product, wu2016aspect]. For instance, to tackle the weakness of the “bag of words” assumption, he2017unsupervised proposed an unsupervised neural network model. Considering dependencies between adjacent words, the authors used an embedding method with attention mechanism to de-emphasize saliency and extract aspects. angelidis2018summarizing proposed a weakly supervised neural framework for the identification and extraction of salient customer opinions that combines aspect and sentiment information. Using a small number of annotated instances with a large-scale unlabeled corpus, bravzinskas2020few proposed a few-shot learning framework for generating an abstractive summary. In recent years, pre-trained language models [vaswani2017attention, kenton2019bert] have been shown to be effective in document summarization [liu2019hierarchical]. A domain-specific generative pre-training method, PEGASUS, was proposed to address the e-commerce review summarization problem [3404835zhang3463037]. Inspired by Vector-Quantized Variational Auto-encoders (VQ-VAE) [van2017neural], angelidis2021extractive proposed an unsupervised neural model, Quantized Transformer (QT), that uses a clustering interpretation of the quantized space to discover popular opinions among hundreds of reviews. To tackle challenges such as a lack of cross item diversity and consistency, ovedpass21 proposed a method that leverages strong pre-trained language models.
@@
To address these problems, dong2017learning proposed an attention-enhanced attribute-to-sequence model to generate item reviews for given attribute information. The authors proposed an attention-enhanced attribute-to-sequence model that learns to encode attributes into vectors and then uses a recurrent neural networks based on LSTM units to generate reviews by conditioning on the encoding vectors; see Fig. 3.13. The model can be divided into three components: an attribute encoder, a sequence decoder, and an attention mechanism. The authors use a dataset collected from Amazon to verify the effectiveness of the proposed method, especially the attention mechanism in the review generation procedure.
@@
There is increasing attention for combining preference prediction with review generation. As sentiment classification plays an important role in e-commerce review analysis, radford2017learning proposed a representation learning strategy to detect opinions while generating reviews, where a generic sentiment tree bank was applied to represent the sentiment label in user reviews [socher2013recursive]. Many e-commerce sites provide structured information, such as aspect-sentiment scores, i.e., each review text contains sentences describing a number of aspects of the item. Focusing on generating long Chinese reviews from aspect-sentiment scores, zang2017towards proposed end-to-end sequential review generation models (SRGMs). Unlike traditional seq2seq models, SRGMs encode inputs of aspect-sentiment scores using multi-layer perceptrons. sharma2018cyclegen proposed an LSTM-based neural network to generate personalized reviews from multi-faceted factors, i.e., user profiles and item attributes, where an additional loss term is used to ensure consistency of the sentiment rating in the generated review.
ni2017estimating proposed a collaborative-filtering generative concatenative network to jointly optimize item recommendation and generate personalized reviews. To generate personalized high-fidelity reviews, ni2018personalized proposed an encoder-decoder model to leverage both user and item information as well as auxiliary, textual input and aspect-aware knowledge, where an attention fusion layer is introduced to control the influence of various encoders.
@@
Some e-commerce sites have launched an interaction box called tips on their mobile platforms. Fig. 3.14 shows examples of reviews and tips on Yelp. The left column is the review from the user “Monica H.”, and tips from several other users are shown in the right column. Tips are more concise than reviews and can reveal user experience, feelings, and suggestions with only a few words. To generate concise tips from reviews, Li2017 proposed a multi-task learning framework for tip generation and rating prediction. For abstractive tip generation, gated recurrent neural networks are employed to decode user and item latent factors, whereas for rating regression, a multilayer perceptron network is employed to project user latent factors and item latent factors into ratings. A persona-aware tip generation framework has been proposed for personalized tip generation through adversarial variational auto-encoders (aVAE) [li2019persona].

### v3

@@ 3.1.1 Basic concepts
An interface refers to an interactive component of a webpage or an application [hearst-2009-search]. Referring to all interactive components on e-commerce portals (e.g., search bars, navigation panels, lists of recommended items, item titles, and user reviews), e-commerce interfaces play an invaluable role for the e-commerce user experience. E-commerce interfaces dramatically impact the performance of an e-commerce platform. Depending on the nature of the stakeholders involved, most e-commerce sites can be divided into four types of business: business-to-business, business-to-consumer, consumer-to-consumer, and consumer-to-business [nemat2011taking]:
@@
B2C refers to scenarios where businesses directly sell items to consumers. Most online shopping platforms, such as Amazon, Booking.com, and JD.com belong to this type of business. Fig. 3.1(2) shows an item page from Amazon as an example of the interface used by B2C businesses.
@@ 3.1.2.1 Search results in e-commerce
E-commerce search engines provide category options with the search bar. During the early development of e-commerce search, interfaces of different types have been considered, e.g., devoted type, divided type, co-existing type, and multi-page type [lu2006clustering]. But with the development of e-commerce search, these types of interfaces have been blended by e-commerce platforms. Currently, a typical e-commerce search system includes three main components: query processing, candidate retrieval and ranking [zhang2020towards]. In query processing, the search engine rewrites a query from the user into a term-based representation that can be processed by downstream components. In the candidate retrieval stage, the system uses the inverted index to retrieve candidate products to match queries. Finally, the ranking component orders the retrieved candidates based on factors such as relevance, and predicted conversion ratio. We will discuss research into the principles and strategies of all three components in Section 5 in more detail.
@@ 3.1.2.2 Recommendation results in e-commerce platforms
For many e-commerce platforms, recommendations have become the most important service to help users find their needed items. E.g., recommendations have been reported to contribute to the majority of both revenue and traffic in Taobao [wang2018billion], where one billion users can be connected to two billion items. To this end, the homepage on the mobile Taobao app is generated based on consumers’ past behavior via recommendation algorithms, as illustrated in Fig. 3.3.
@@
Different types of recommendation results may be shown at different stages of a customer’s. Examples include “substitutes” (see Fig. 3.4(a) and “complementary items” before and after the user adds a product to their cart (see Fig. 3.4(b) and 3.4(c), respectively). Once the consumer clicks a recommended product, the system will automatically jump to the product detail page, which includes product titles, product descriptions, categories, ratings, and reviews. We will discuss more details about strategies and technologies of e-commerce recommendation in Section 6.
@@ 3.1.2.3 Product titles in e-commerce platforms
Item title compression, also called short title extraction [gong2019automatic], is meant to extract sufficient words from lengthy and verbose titles to produce a succinct new title to improve the user experience on mobile devices [wang2018multi]. Inspired by neural extractive document summarization methods [ren2017leveraging], item title compression methods apply neural networks to weight the importance of each word in the item title. gong2019automatic introduce a feature-enriched neural extractive model to extract short titles. Specifically, the authors apply a recurrent neural network as a sequential classifier with three types of features: content, attention, and semantics respectively. By using user search logs as external knowledge, wang2018multi construct a multi-task learning approach for improving item title compression. The proposed method is composed of two seq2seq components which share an identical encoder. The authors combine these two components with an overall pointer neural network [vinyals2015pointer] to automatically select the most informative words from the given item title.
Pointer neural networks easily omit key information. To tackle this problem, sun2018multi introduce a multi-source pointer network model, named the multi-source pointer network (MS-Pointer), by considering two extra constraints: (i) irrelevant information reduction; and (ii) the key information retainment. Fig. 3.6 provides an overview of MS-Pointer, with two encoders. In MS-Pointer, in addition to the encoder for the source title, the authors add another knowledge encoder that uses an LSTM to embed the brand name and the commodity name. As shown in Fig. 3.6, MS-Pointer combines the original title “Nintendo switch console…” and background knowledge “brand name: Nintendo”, and then it generates the short title about the item “Nintendo switch”. More recently, DBLP:conf/emnlp/FetahuCRM23 have proposed an instruction fine-tuning strategy to summarize product titles according to various criteria such as the number of words in a summary or the inclusion of specific phrases.
@@
The task of title generation has been proposed to extend the task of title compression into a text generation problem. Unlike title compression, which only extracts words from item titles, the task of title generation is to generate a short item title so as to address the problem of inaccurate item titles in e-commerce [zhang2019multi]. To generate a succinct and accurate short title from a long source title, zhang2019multi offer a multi-modal generative adversarial network, named MM-GAN, which addresses the title generation task as a reinforcement learning problem. MM-GAN is composed of two main components, a title generator and a discriminator (Fig. 3.7). Given the source title and its corresponding tags or features, the generator applies an LSTM-based network to generate a short item title. The discriminator, i.e., a binary classifier, distinguishes whether the generated short titles are human-generated or machine-generated. Thus, an adversarial learning procedure is constructed, in which the quality of the short title depends on its ability to fool the discriminator into believing it is a human-generated one, and the output of the discriminator is a reward for the generator to improve the generation performance. Recently, scene marketing has become a new marketing mode for product promotion where scene scenarios are created to demonstrate product functions [zhao2020data]. To help the e-commerce system find scene topics, lin2022automatic propose a topic generation method to generate scene-based titles in e-commerce.
@@ 3.1.2.4 Product descriptions in e-commerce platforms
As shown in Fig. 3.8, many e-commerce platforms provide a short description for each item so as to showcase the features of the item. As an important factor in content marketing, the item description is key for increasing consumer engagement. During the early years of e-commerce, item descriptions were usually written or edited by human copywriters. However, the availability of an increasing number of items in e-commerce makes this manual process too costly. Moreover, with the development of virtual assistants in e-commerce, such as Alexa and Tmall Genie, there is a growing demand for automatically generating a short description given item attributes. To address this demand, the task of item description generation has been proposed. Item description generation needs to generate an item’s description from a series of complicated attributes. wang2017statistical detail a statistical framework to weight the relative importance of the attributes of an item and to maintain accuracy at the same time. In Fig. 3.9 we specify the framework of the proposed item description model. By combining sentence-level templates extracted from the input data with knowledge from a pre-trained dataset, the authors generate and rank candidate item descriptions through an online document planning stage.
@@
Unlike early studies that focused on generating item descriptions purely from the item’s attributes, zhang2019automatic generate a pattern-controlled item description from multiple features, e.g., titles and item categories. Based on the copy mechanism [gu2016incorporating], the authors propose a pattern-controlled pointer-generator network (PGPCN) to generate the description. In PGPCN, a transformer is applied in the encoder component, whereas the decoder is used to control the pattern (e.g., the category, the length, and the style of the description) of the item.
It is important that the descriptions generated for item description are grounded in facts. To generate a fact-based description, chanstick2019 offer an encoder-decoder framework, called the fidelity-oriented product description generator (FPDG), by searching key information from keyword labels. The authors establish semantic connections between item keywords and the generated product description. As shown in Fig. 3.10, FPDG has two main components: (i) a keyword encoder that stores the word and its entity label in the token memory and self-attention modules, and (ii) an entity-based generator that generates an item description based on the memory and self-attention modules.
@@
Personal interest is neglected by all of the above approaches that generate descriptions given attributes or keywords. To address this shortcoming, chen2019towards propose a knowledge-based personalized item description generation strategy. The authors extend the encoder-decoder framework [sutskever2014sequence] to a sequence modeling formulation using a self-attention mechanism. A large variety of item attributes, including the target user’s personalized preference features, are combined in an attribute fusion component through multi-layer attention mechanisms; retrieved external knowledge is incorporated in a knowledge incorporation component. In Fig. 3.11, we provide an example of the knowledge-based personalized item description generation.
@@ 3.1.2.5 Question answering in e-commerce
Question-answering (QA) systems are designed to provide direct and concise answers to user queries by understanding the intent behind a question and retrieving or generating the most relevant information. QA systems aim to deliver specific answers from structured or unstructured data sources, such as web documents or knowledge bases. QA systems are essential across a variety of domains, including web search, customer support, where users seek quick, accurate, and contextually relevant information [radev2002evaluating, tapeh2008knowledge, yin2015neural]. To increase the number of sales, most e-commerce portals provide a QA service to facilitate the customers’ shopping procedure by answering their questions about products [gao2019product, feng2021multi]. Currently, on many e-commerce sites, a user can ask a question about a product, and the QA system allows some users (e.g., customers who bought this product) to provide answers [gao2019product]. In Fig. 3.12, we show examples of question-answering services at Amazon and JD.com. More detailed discussions of e-commerce QA are provided in Section 7.1.
@@ 3.1.2.6 User reviews in e-commerce
The sentiment classification task is to label a given text with a specific opinion label. It has received lots of attention during the past two decades [pang2002thumbs, go2009twitter, pan2010cross, kamal2012mining, tang2014learning, pontiki2015semeval, tsytsarau2016managing, sun2017review]. Work on sentiment classification in an e-commerce context has attempted to capture a user’s opinion about a specific item from reviews on an e-commerce platform [sun2017review, tang2015learning, xia2015dual, tripathy2016classification, li2018effect]. Traditional approaches to sentiment classification focus on the classification problem given textual attributes of an item [pang2002thumbs], while largely ignoring the relation between users and the item. To address this problem, tang2015learning introduce a neural network, the user-product neural network (UPNN), to incorporate user and item information into a document-level sentiment classification procedure. In particular, the authors jointly embed the user preference information, i.e., ratings and reviews, and item attributes. Then, a convolutional neural network is used to predict the sentiment label of the target review.
Sentiment classification using the relation between users and products or services faces two important challenges: (i) the sparseness of user-item interactions, and (ii) the information in user embedding methods. chen2016neural present a fine-grained hierarchical neural network model to incorporate global user and item information into sentiment classification. Unlike many sentiment classifiers that use convolutional neural networks, the authors apply a hierarchical LSTM [gers1999learning] to jointly generate sentence-level representations and document-level representations. Then, user-item interaction information is applied as attention over various regions of a document to enhance the sentiment classification. wu2018improving distinguish between different roles of words and sentences in user reviews: to describe the user’s preferences or to describe an item’s characteristics. To distinguish between these roles, the authors put forward an attention-based hierarchical neural network model to embed user and item information to generate two text representations with user attention or item attention, respectively. Hao2021wwwsentiment use fine-grained latent opinion knowledge into the sentiment classification process by using a variational reasoning method.
@@
Given the fact that an item can be commented on by hundreds of thousands of consumers, the quality of reviews in e-commerce varies considerably and not all reviews are helpful. To gain insights from helpful reviews, the task of review helpfulness prediction has attracted attention from both academia and industry [mcauley2016addressing]. Early studies on review helpfulness prediction employ feature-aware methods, where multiple types of features, such as structural features [kim2006automatically, susan2010makes, xiong2011automatically], emotional features [martin2014prediction], semantic features  [yang2015semantic], argument features [liu2017using], and lexical features [xiong2014empirical], are successfully applied.
Motivated by the progress of deep neural networks, fan2018multi introduce a multi-task neural learning (MTNL) architecture for identifying helpful reviews. chen2018cross propose a CNN-based neural network with multi-granularity (i.e., character-level, word-level, and topic-level) features for helpfulness prediction. fan2019product suggest an end-to-end deep neural architecture to capture the intrinsic relationship between the meta-data of an item and its numerous comments that could be beneficial to discover the helpful reviews. Multi-modal data has become increasingly popular in online reviews. To analyze multi-modal reviews, liumulti2021 introduce a multi-modal review helpfulness prediction task that is aimed at exploring multi-modal clues for review helpfulness prediction. The authors describe an item-review coherent reasoning module to capture the intra- and inter-modal coherence between the target item and the review. han2022sancl put forward a selective attention approach, including probe mask generation and mask-based attention computation, for the multi-modal review helpfulness prediction problem. To mine the mutual information of cross-modal relations in the input, nguyen2022adaptive propose an adaptive cross-modal contrastive learning mechanism, with a multi-modal interaction module to correlate modalities’ features.
@@
Feature-aware methods are inspired by previous document summarization methods: yang2010feature detail a feature-based item review summarization method to satisfy the detailed information needs of customers. To jointly summarize reviews and predict ratings in a mobile environment, liu2012movie offer a latent semantic indexing based approach to extract features and attributes from user reviews and ratings. For new items without reviews, a probabilistic retrieval method is proposed to extract relevant opinion features from other items to describe the item information [park2015retrieval]. Another important part of review summarization concerns aspect extraction from user reviews [chen2014aspect, angelidis2018summarizing, bravzinskas2020few], where target entities and aspects need to be extracted from opinionated text. chen2014aspect extract prior knowledge automatically from user reviews and propose a fault-tolerant model to extract aspects guided by the knowledge. Category hierarchy information is combined with a topic model to improve the performance of aspect extraction [yang2017aspect]. By jointly considering fine-grained aspect-topic-sentiment connections, tan2017sentence propose a generative topic aspect sentiment model.
With the development of deep learning, item review summarization has been tackled from a range of perspectives [ly2011product, wu2016aspect]. For instance, to tackle the weakness of the “bag of words” assumption, he2017unsupervised propose an unsupervised neural network model. Considering dependencies between adjacent words, the authors used an embedding method with attention mechanism to de-emphasize saliency and extract aspects. angelidis2018summarizing describe a weakly supervised neural framework for the identification and extraction of salient customer opinions that combines aspect and sentiment information. Using a small number of annotated instances with a large-scale unlabeled corpus, bravzinskas2020few suggest a few-shot learning framework for generating an abstractive summary. In recent years, pre-trained language models [vaswani2017attention, kenton2019bert] have been shown to be effective in document summarization [liu2019hierarchical]. A domain-specific generative pre-training method, PEGASUS, has been proposed to address the e-commerce review summarization problem [3404835zhang3463037]. Inspired by Vector-Quantized Variational Auto-encoders (VQ-VAE) [van2017neural], angelidis2021extractive explain an unsupervised neural model, Quantized Transformer (QT), that uses a clustering interpretation of the quantized space to discover popular opinions among hundreds of reviews. To tackle challenges such as a lack of cross item diversity and consistency, ovedpass21 offer a method that uses strong pre-trained language models.
@@
To address these problems, dong2017learning propose an attention-enhanced attribute-to-sequence model to generate item reviews for given attribute information. The authors introduce an attention-enhanced attribute-to-sequence model that learns to encode attributes into vectors and then uses a recurrent neural networks based on LSTM units to generate reviews by conditioning on the encoding vectors; see Fig. 3.13. The model can be divided into three components: an attribute encoder, a sequence decoder, and an attention mechanism. The authors use a dataset collected from Amazon to verify the effectiveness of the proposed method, especially the attention mechanism in the review generation procedure.
@@
There is increasing attention for combining preference prediction with review generation. As sentiment classification plays an important role in e-commerce review analysis, radford2017learning describe a representation learning strategy to detect opinions while generating reviews, where a generic sentiment tree bank was applied to represent the sentiment label in user reviews [socher2013recursive]. Many e-commerce sites provide structured information, such as aspect-sentiment scores, i.e., each review text contains sentences describing a number of aspects of the item. Focusing on generating long Chinese reviews from aspect-sentiment scores, zang2017towards offer end-to-end sequential review generation models (SRGMs). Unlike traditional seq2seq models, SRGMs encode inputs of aspect-sentiment scores using multi-layer perceptrons. sharma2018cyclegen propose an LSTM-based neural network to generate personalized reviews from multi-faceted factors, i.e., user profiles and item attributes, where an additional loss term is used to ensure consistency of the sentiment rating in the generated review.
ni2017estimating put forward a collaborative-filtering generative concatenative network to jointly optimize item recommendation and generate personalized reviews. To generate personalized high-fidelity reviews, ni2018personalized come up with an encoder-decoder model to use both user and item information as well as auxiliary, textual input and aspect-aware knowledge, where an attention fusion layer is introduced to control the influence of various encoders.
@@
Some e-commerce sites have launched an interaction box called tips on their mobile platforms. Fig. 3.14 shows examples of reviews and tips on Yelp. The left column is the review from the user “Monica H.”, and tips from several other users are shown in the right column. Tips are more concise than reviews and can reveal user experience, feelings, and suggestions with only a few words. To generate concise tips from reviews, Li2017 suggest a multi-task learning framework for tip generation and rating prediction. For abstractive tip generation, gated recurrent neural networks are employed to decode user and item latent factors, whereas for rating regression, a multilayer perceptron network is employed to project user latent factors and item latent factors into ratings. A persona-aware tip generation framework has been put forward for personalized tip generation through adversarial variational auto-encoders (aVAE) [li2019persona].